# Deep Reinforcement Learning: A Convex Optimization Approach

### v3

@@ 4 Proof of Theorem 1
$\Delta_{K}\leq\prod_{k=1}^{K}(1-\mu\alpha_{k})\Delta_{0}+\sum_{k=1}^{K}\prod_{i=k+1}^{K}(1-\mu\alpha_{i})\alpha_{k}\frac{\rho f_{\textup{max}}}{\lambda T}\|w^{\star}\|$ (11)
@@
$\lim_{K\rightarrow\infty}\frac{e^{\mu}}{\mu}\frac{1}{K^{\mu}}\left((K+1)^{\mu}-1\right)=\frac{e^{\mu}}{\mu}$
@@
$\sum_{k=1}^{\infty}\prod_{i=k+1}^{\infty}(1-\mu\alpha_{i})\alpha_{k}$ $\leq\frac{e^{\mu}}{\mu}$ (15)
@@
$\|w-w^{\star}\|=\Delta_{\infty}\leq C\cdot\frac{\rho}{T}$
@@ 5 Conclusions
$\|w_{k}-w^{\star}\|\leq C\cdot\frac{\rho}{T},$
for some constant $C$. In particular, we can see that as regularization parameter $\rho$ decreases and as the time horizon $T$ increases, the converging neural network parameters get arbitrarily close to the optimal ones. Future work includes applications to reinforcement learning for mixed continuous and discrete state and action spaces. It would also be interesting to apply the algorithm to fine-tuning Large Language Models, since the training would be fast and computationally very efficient, and convergence to the optimal neural network parameters is guaranteed.

### v5

@@ 4 Proof of Theorem 1
$\Delta_{K}\leq\prod_{k=1}^{K}(1-\mu\alpha_{k})\Delta_{0}+\sum_{k=1}^{K}\prod_{i=k+1}^{K}(1-\mu\alpha_{i})\alpha_{k}\frac{\rho f_{\textup{max}}}{\lambda}\|w^{\star}\|$ (13)
@@
$\lim_{K\rightarrow\infty}\frac{e^{\mu}}{\mu}\frac{1}{K^{\mu}}\left((K+1)^{\mu}-1\right)=\frac{e^{\mu}}{\mu},$
@@
$\sum_{k=1}^{\infty}\prod_{i=k+1}^{\infty}(1-\mu\alpha_{i})\alpha_{k}$ $\leq\frac{e^{\mu}}{\mu}.$ (17)
@@
$\|w-w^{\star}\|=\Delta_{\infty}\leq C\cdot{\rho}$
@@ 5 Conclusions
$\|w-w^{\star}\|\leq C{\rho},$
for some constant $C$. In particular, we can see that as regularization parameter $\rho$ decreases, the converging neural network parameters get arbitrarily close to the optimal ones. As a consequence, our algorithm converges fast due to the polynomial-time convergence of convex optimization algorithms.
Future work includes applications to reinforcement learning for mixed continuous and discrete state and action spaces. It would also be interesting to apply the algorithm to fine-tuning Large Language Models, since the training would be fast and computationally very efficient, and convergence to the optimal neural network parameters is guaranteed.